The intensity in I = (1/2?0)EmaxBmax and in  I = E2max/(2?0c) = (cB2max)/(2?0) is written in terms of the maximum electric and magnetic fields. Rewrite these equations in terms of the rms field values. (Use the following as necessary: Brms, Erms, c, and ?0. Do not substitute numerical values; use variables only.)

a) I = (1/2?0)EmaxBmax =---------------------------

b) I = E2max/(2?0c) =______________________

c)   I  = (cB2max)/(2?0)=_____________________
Answer #1

Erms = Emax/sqrt 2

Emax = sqrt 2*Erms

Similarly

Bmax = sqrt 2*Brms

Part A

I = \frac{1}{2\mu _{0}}E_{max}B_{max} = \frac{1}{2\mu _{0}}\sqrt{2}E_{rms}*\sqrt{2}B_{rms}

I = \frac{1}{2\mu _{0}}2E_{rms}B_{rms} = \frac{E_{rms}B_{rms}}{\mu _{0}}

Part B

I = \frac{E_{max}^2}{2\mu _{0}c} = \frac{(\sqrt 2E_{rms})^2}{2\mu _{0}c}

I = \frac{2E_{rms}^2}{2\mu _{0}c} = \frac{E_{rms}^2}{\mu _{0}c}

Part C

I = \frac{cB_{max}^2}{2\mu _{0}} = \frac{c(\sqrt 2E_{rms})^2}{2\mu _{0}}

I = \frac{2cB_{rms}^2}{2\mu _{0}} = \frac{cB_{rms}^2}{\mu _{0}}

Please Upvote.